Aggieland Enterprise just paid $400,000 to buy ten new company trucks for its salesperson to use. Aggieland expects to use these cars for five years and then replace them. If the salvage value of 5 of the cars is $2,000 each, and for the rest of the cars it is $10,000 each. How much depreciation expense will Aggieland record each year using the straight-line method
Answer #1

Answer : $ 68,000

Annual Depreciation = (cost -Salvage Value )/ Life in Years

Cost = 400,000

Salvage Value =(5 *2,000)+(5*10,000) =60,000

Annual Depreciation = (400,000-60,000)/ 5 = $ 68,000 (Answer)